two trains left the same station at the same time and traveled in opposite directions. the E train avereged 130km/hr. The A train's speed was 110km/h. In how many hours were hey 480km apart?
Answer by checkley75(3666)

130T+110t=480
240T=480
T=480/240
T=2 HOURS THEY WILL BE 480km APART.
PROOF
130*2+110*2=480
260+220=480
480=480